# Parabolic bundles and the intersection cohomology of moduli spaces of vector bundles on curves

### v1

@@ 4.1. Towers of projective spaces
The space $\mathcal{T}_{(V_{1},\ldots,V_{k})}$ defined above is a smooth projective variety of dimension $\sum_{i<j}r_{i}r_{j}(g-1)+r-k$, which is an iterated fibration by projective spaces.
Remark 4.2.
@@
$\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})=\{\varphi:\mathcal{W}^{\underline{i}}\to\zeta^{*}V_{i+1}\;|\\ \forall w\in{\mathcal{T}}_{(V_{1},\ldots,V_{i})}\,\,\,\varphi|_{\{w\}\times C}\text{\, is a morphism of parabolic bundles}\}.$
@@
(iii) The normalized universal bundle on $\mathcal{P}_{0}(r_{1}+\ldots+r_{i+1})\times C$ (cf. notation on page 3.9) restricts to the universal bundle $\mathcal{W}^{\underline{i+1}}$ on $\mathcal{T}_{(V_{1},\ldots,V_{i+1})}\times C$.
@@
To calculate its dimension, observe that for $j=2,\ldots,k$ the fiber dimension of $\psi_{j-1}$ is equal to
@@
To obtain the first equality we applied the long exact sequence (20) to the pair of parabolic bundles $W/W_{j}$ and $V_{j}$; for the last equality we used that $\mathrm{parslope}(W/W_{j})<\mathrm{parslope}(V_{j})$, and thus the sheaf $\mathrm{ParHom}(W/W_{j},V_{j})$ has no nontrivial global sections (cf. [30, Proposition 6]).
@@ 4.2. The cohomology ring of the projective tower
It is easy to see that $\mathbb{Q}[x_{1},\ldots,x_{k}]/I_{(r_{1},\ldots,r_{k})}$ has the Hilbert function (30), thus it is sufficient to identify the relations $\phi_{i}$, $i=1,\dots,k$, in the cohomology of our tower.
@@ 4.3. The Poincaré polynomial of the abelian fiber
Now we begin the study of the cohomology of the fibers of the map $\pi$. Fix a partition $\rho=[r_{1},r_{2},\dots,r_{k}]\vdash r$. Recall the notation $F_{V}:=\pi^{-1}(V)$ for the fiber of the map $\pi$.
The following statement is straightforward.
@@
    $F_{V}=\bigcup_{\begin{subarray}{c}\sigma\in\Sigma_{k}\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})},$
@@
Our next goal is to provide a recursive formula for the Betti numbers of the abelian fibers. We will need the following statement.
Lemma 4.11.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, $V\in M_{\rho}^{\mathrm{ab}}$. Then
  1. (i)
    the weight polynomial (see §3.5.3) of the fiber $F_{V}$ coincides with its Poincaré polynomial: $E_{t}(F_{V})=P_{t}(F_{V})$, and
  2. (ii)
    the odd cohomologies of $F_{V}$ vanish.
@@
Our starting point is the representation of the cohomology of $F_{\rho}$ via the Decomposition Theorem (see (25)). It is easy to verify that abelian strata are in the closure of abelian strata only. As abelian strata have no local monodromy, Proposition 3.14, and Corollary 3.17 imply that the cohomology of $F_{V}$ is pure. This means, in particular, that that the weight polynomial of $F_{V}$ coincides with its Poincaré polynomial: $E_{t}(F_{V})=P_{t}(F_{V})$.
According to Proposition 4.10, we have a surjective birational map from a disjoint union of smooth towers to $F_{V}$:
$\sqcup_{\sigma\in\Sigma_{k}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})}\to F_{V}.$
Now using Proposition 3.13, we can conclude that the map
(35) $H^{\bullet}(F_{V})\hookrightarrow\oplus_{\sigma\in\Sigma_{k}}H(\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})})$
is an injection, and thus (see Proposition 4.5) the fiber $F_{V}$ has only even cohomology. ∎
@@
Note that $\mathcal{F}_{i}$, is the set of those elements of the fiber $F_{V}$ which have $V_{i}$ as a subbundle. It follows then, that for a subset $J\subset\underline{k}$, the set
@@
Since $\mathcal{F}_{i}$ is a projective bundle (37) with fiber dimension $r_{i}(r-r_{i})(g-1)+r_{i}-1$ over $F_{V/V_{i}}$, its cohomology is pure (cf. Lemma 4.11(i)), and thus we can identify the term in (36) corresponding to $J=\{i\}$ as the weight polynomial of $\mathcal{F}_{i}$.
@@
Now observe that $\{\mathcal{F}_{i},\,i=1,\dots,k\}$ is a covering of $F_{V}$, and the corresponding inclusion-exclusion formula results in the identity
@@
This is exactly the equality (36) (cf. Lemma 4.11), and this completes the proof of Theorem 4.12. ∎
@@ 4.4. Intersection of projective towers
Our next project is to study the intersection between two towers, which will be key in §7. Our construction is a generalization of the tower construction of §4.1. To describe this generalization, we will need to introduce a combinatorial device associated to a pair of permutations: $\sigma,\tau\in\Sigma_{k}$.
@@
Using this partial order, we construct and ordered decomposition of the set $\underline{k}$
@@
Clearly, this process eventually exhausts all of $\underline{k}$.
@@
    $\mathcal{I}(\sigma,\tau)^{\leq{1}}=\prod_{i\in\Pi_{1}}\mathbb{P}(V_{i|p}),$
@@
    together with the normalized universal bundle (cf. notation on page 3.9 and Lemma 4.3(iii))
@@
Clearly, this happens only if $j<i$. In other words, for any $W\in\mathcal{T}_{\sigma}$ we obtain that $j\prec_{W}i$ implies $\sigma^{-1}(j)<\sigma^{-1}(i)$. Hence if $W\in\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}$ then the partial order $\prec_{W}$ has to be compatible with the partial order
@@ 4.5. The cohomology ring of the abelian fiber
Our goal is to show that $\phi$ is an isomorphism. In order to show this, we use the following statement, which we will prove in §6.4.
@@
To prove the statement, we need to show that $\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\subset I$. Let $d(\rho)=\sum_{1\leq i<j\leq k}r_{i}r_{j}(g-1)$; we start with proving that the surjective morphism of graded rings $\eta$:
@@
induces an isomorphism between the homogeneous degree $2d(\rho)+2r-2k$ elements. It follows from Proposition 4.17 and the recursive formula (36), that the ring $R$ has $k!$ elements of top degree $2d(\rho)+2r-2k$. Since the top-dimensional cohomology of an equidimensional union of projective varieties equals the direct sum of the pieces, we have (see Propositions 4.1 and 4.10(i))
@@
Then using Lemma 4.15 and the integral formula (34), we obtain that for every $\sigma\in\Sigma_{k}$ the class
@@
under the $\sigma$-th coordinate projection in the direct sum (42). Since the morphism $\theta$ is injective (see the proof of Lemma 4.11), and the $k!$-dimensional space
@@
Next, we repeat the argument from [17, proof of Theorem 7.1]: we note that the ring $R$ can be interpreted as the cohomology ring of the Lawrence toric variety of dimension $d(\rho)+r-k$ (cf. [31, Theorem 6.3] and [17, Theorem 8.3]), and thus the socle of $R$ consists precisely of the elements of top degree $2d(\rho)+2r-2k$. We assume that $\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\not\subset I$ and pick a non-zero element $h\in\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\subset R$ of maximal degree. Our goal is to show that $h$ is zero in $R$, and thus $h\in I$.
@@ 5. The no-support theorem for non-abelian strata
Finally, we turn to the general case of the fibers. Recall from §3.1 that a complete stratification $\mathcal{M}_{0}(r)=\sqcup_{\overline{\rho}}M_{\overline{\rho}}$ is parametrized by multi-partitions:
@@
A non-abelian stratum $M_{\overline{\rho}}$ carries no support of $R\pi_{*}\mathbb{Q}_{\mathcal{P}_{0}(r)}$, thus the corresponding local system $\mathcal{L}_{\overline{\rho}}$ vanishes.
@@
In fact, we will prove a stronger statement. Recall from §3.1 that to the multi-partition $\overline{\rho}$ corresponds a true partition $\rho$ with $k=m_{1}+\dots+m_{s}$ elements. Now let $V\in M_{\rho}^{\mathrm{ab}}\cap\mathcal{C}^{\circ}$; we will show that the natural restriction map
$g:H^{\bullet}(F_{\overline{V}})\longrightarrow H^{\bullet}(F_{V})$
is injective, which clearly implies the injectivity of $j^{*}$.
To the multi-partition ${\overline{\rho}}$, we associate the subgroup $\Sigma[\overline{\rho}]\simeq\Sigma_{m_{1}}\times\dots\times\Sigma_{m_{s}}\subset\Sigma_{k}$, which exchanges the elements of ${\underline{k}}=\{1,2,\dots k\}$. The group $\Sigma[\overline{\rho}]$ acts naturally on the cohomology of an abelian fiber $F_{V}$, and, as we will show in Proposition 5.2, the image of the restriction map $g$ is contained in $H^{\bullet}(F_{V})^{\Sigma[\overline{\rho}]}$.
5.1. Injectivity and the cohomological map
The final statement is as follows.
@@ 5.1. Injectivity and the cohomological map
Let $\overline{\rho}$ be a multi-partition and $\rho$ the associated partition, $\overline{V}\in M_{\overline{\rho}}$, and let $V\in M_{\rho}^{\mathrm{ab}}\cap U$ be as above. The map
@@
To simplify our notation, we will describe in detail the case of $\overline{\rho}=[(s,k)]$, that is, the stratum $M_{\overline{\rho}}$ of bundles $\overline{V}=W^{\oplus k}$, where $W$ is a stable bundle of rank $s$ and degree 0. The general case is obtained by tensoring copies of this basic example (see Remark 5.12).
We thus have $sk=r$, and $\rho=[s,s,\dots,s]$. To avoid double indices, we introduce the notation $lW=W^{\oplus l}$. We note that the stratum $M_{\rho}^{\mathrm{ab}}$ may be obtained as a quotient of the space of sequences of stable bundles $\widetilde{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},\dots,V_{k})|\,V_{1}\oplus...\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}\}$ by the symmetric group $\Sigma_{k}$. Then the cohomologies of the abelian fibers $F_{\oplus V_{i}}$ form a trivial bundle over $\widetilde{M}_{\rho}^{\mathrm{ab}}$, to which the action of $\Sigma_{k}$ lifts as the permutation action on the variables $x_{j}$, $j=1,\dots,k$. It follows then that the image of the composed map (43) is contained in the invariant part of the cohomology of $H^{\bullet}(F_{V})$; we will show now that the map
@@
is an isomorphism. Using induction, we assume that the statement holds for $k$ replaced by $k-1$, and our goal is to prove it for $k$. Just as in the abelian case, we begin with establishing the purity of the cohomology of the fiber.
Lemma 5.3.
The cohomology of $F_{kW}$ is pure.
Proof.
The argument is similar to the one given in the abelian case in the proof of Lemma 4.11. Our starting point is (23), which expresses the cohomology of $F_{kW}$ as a direct sum of contributitions from lower strata. Using an inductive hypothesis, we can assume that the lower non-abelian strata carry no support, and thus we can ignore them from the point of view of the Decomposition Theorem. This, up to taking products of our example, reduces the question of purity to the study of the contribution of the stratum $M_{\rho}^{\mathrm{ab}}$ with $\rho=[s,s,\dots,s]$. In this case, the monodromy group is $\Sigma_{k}$, and the variation of the Hodge structure of $H^{\bullet}(F_{V})$ as $V$ varies is trivial, since $H^{\bullet}(F_{V})$ has cohomology of Hodge-Tate type only (see (35)). Now, again we appeal to Proposition 3.14, and Corollary 3.17 to conclude that the cohomology of $F_{kW}$ is pure. ∎
@@
where the map $f^{\prime}$ is a smooth fibration. In fact, we will be mainly interested by the subset ^$\hat{\mathcal{E}}_{\rho}$ of $\hat{\mathcal{E}}$ lying over the closure of $M_{\rho}^{\mathrm{ab}}\subset\mathcal{M}_{0}(ks)$.
We introduce the notation
@@
for the composed map $\widehat{\pi}=\pi\circ h^{\prime}$. It follows from the proof of Proposition 4.1 that for an abelian $V=V_{1}\oplus\dots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ we have
@@
The purity of $H^{\bullet}(F_{V})$ (see Lemma 4.11) and that of the cohomology of the sets $\mathcal{F}_{j}$ (see (37)), which form a covering of $F_{V}$, imply that we have a $\Sigma_{k}$-invariant embedding $H^{\bullet}(F_{V})\hookrightarrow\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$, and this provides us with an embedding
$Q:H^{\bullet}(F_{V})^{\Sigma_{k}}\hookrightarrow\left[\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})\right]^{\Sigma_{k}}.$
@@
By Lemma 5.3, $H^{\bullet}(F_{kW})$ has pure Hodge structure, and so does $H^{\bullet}(\mathcal{E})$, since $\mathcal{E}$ is a bundle of projective spaces over $F_{(k-1)W}$. Now the injectivity follows from Proposition 3.13. ∎
Finally, we note that the cohomology $H^{\bullet}(\widehat{\pi}^{-1}(V))$ of the fibers (46) over $M_{\rho}^{\mathrm{ab}}$ may be obtained as a descent to $M_{\rho}^{\mathrm{ab}}$ of a trivial bundle $\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$ on $\widetilde{M}_{\rho}^{\mathrm{ab}}$ endowed with a $\Sigma_{k}$-action. This results in a natural map
@@
Using descent with respect to the subgroup $\Sigma_{k-1}\subset\Sigma_{k}$, we can conclude that the projection of the map $H^{\bullet}(\mathcal{E})\longrightarrow\bigoplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$ to $H^{\bullet}(\mathcal{F}_{1})$ has its image in $H^{\bullet}(\mathcal{F}_{1})^{\Sigma_{k-1}}$, and recall that by our inductive hypothesis
@@
Clearly, Lemma 5.5, and the commutative diagram (48) imply that the map (43) is injective. This already proves Theorem 5.1. For completeness, however, we will show below that this map is an isomorphism. ∎

### v2

@@ 5.1. Towers of projective spaces
The space $\mathcal{T}_{(V_{1},\ldots,V_{k})}$ is an iterated fibration by projective spaces of dimension $\sum_{i<j}r_{i}r_{j}(g-1)+r-k$.
@@
$\mathrm{ParHom}(\mathcal{W}^{\underline{i}},\zeta^{*}V_{i+1})=\{\varphi:\mathcal{W}^{\underline{i}}\to\zeta^{*}V_{i+1}\;|\\ \forall w\in{\mathcal{T}}_{(V_{1},\ldots,V_{i})}\,\,\,\varphi|_{\{w\}\times C}\text{\, is a morphism of parabolic bundles}\}$
@@
(iii) The normalized universal bundle on $\mathcal{P}_{0}(r_{1}+\ldots+r_{i+1})\times C$ (cf. notation on page 3.12) restricts to the universal bundle $\mathcal{W}^{\underline{i+1}}$ on $\mathcal{T}_{(V_{1},\ldots,V_{i+1})}\times C$.
@@
To calculate its dimension, observe that for $j=2,\ldots,k$, the fiber dimension of $\psi_{j-1}$ is equal to
@@
To obtain the first equality, we applied the long exact sequence (20) to the pair of parabolic bundles $W/W_{j}$ and $V_{j}$; for the last equality we used that $\mathrm{parslope}(W/W_{j})<\mathrm{parslope}(V_{j})$, and thus the sheaf $\mathrm{ParHom}(W/W_{j},V_{j})$ has no nontrivial global sections (cf. [34, Proposition 6]).
@@ 5.2. The cohomology ring of the projective tower
It is easy to see that $\mathbb{Q}[x_{1},\ldots,x_{k}]/I_{(r_{1},\ldots,r_{k})}$ has Hilbert function (30), thus it is sufficient to identify the relations $\phi_{i}$, $i=1,\dots,k$, in the cohomology of our tower.
@@ 5.3. The Poincaré polynomial of the abelian fiber
Now we begin the study of the cohomology of the fibers of the map $\pi$. The following statement is straightforward.
Fix a partition $\rho=[r_{1},r_{2},\dots,r_{k}]\vdash r$ and recall the notation $F_{V}:=\pi^{-1}(V)$ for the fiber of the map $\pi$.
@@
    (35) $F_{V}=\bigcup_{\begin{subarray}{c}\sigma\in\Sigma_{k}\end{subarray}}\mathcal{T}_{(V_{\sigma(1)},\ldots,V_{\sigma(k)})},$
@@
We first prove the following important purity result.
Proposition 5.10.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$ and $V\in M_{\rho}^{\mathrm{ab}}$. Then the MHS of the cohomology $H^{\bullet}(F_{V})$ of the fiber at $V$ is pure.
@@
We want to prove the result by applying Proposition 4.6 (1), thus we need verify that its hypotheses are satisfied by the map $\pi$ in a neighborhood of the abelian bundle $V$. First, note that abelian strata are in the closure of abelian strata only, so that the the strata appearing equation 25 are indexed by partitions.
We proceed by induction on the length of the partition $\rho$: if $\rho$ has length 1 then $F_{V}$ is isomorphic to $\mathbb{P}^{r-1}$, and purity is clear. Suppose now that $\rho$ is a refinement of length $l$ partition $\mu$. By the inductive hypothesis, MHS of the fiber cohomology $H^{\bullet}(F_{W})$ is pure for $W\in M_{\mu}^{\mathrm{ab}}$, and thus, by Proposition 5.4, the map (see (35))
(36) $H^{d}(F_{W})\hookrightarrow\oplus_{\sigma\in\Sigma_{l}}H^{d}(\mathcal{T}_{(W_{\sigma(1)},\ldots,W_{\sigma(l)})})$
is an injection respecting Hodge structures. The direct sum of the cohomologies of the smooth towers in (36) clearly satisfies the following properties:
  • the cohomology groups vanish in odd degrees;
  • they form a constant system in $D_{V}\cap M_{\mu}^{\mathrm{ab}}$, where $D_{V}$ is a small neighborhood of $V$;
  • this system underlies a trivial variation of Hodge structure of degree $d$.
Denote by $\mathcal{H}^{d}(F_{\mu})$ the local system on $M_{\mu}^{\mathrm{ab}}$ given by the cohomology of the fiber. The injection of MHS in (36) implies that $\mathcal{H}^{d}(F_{\mu})$ enjoys the three properties above on $D_{V}\cap M_{\mu}^{\mathrm{ab}}$ and so does the local system $\mathcal{L}_{\mu}^{d}\subset\mathcal{H}^{d}(F_{\mu})$. Thus our map $\pi$ satisfies conditions (a) and (b) of Proposition 4.6 (1); condition (c) is satisfied by Corollary 4.3, and we can conclude that $H^{\bullet}(F_{V})$ has pure Hodge structure. ∎
Remark 5.11.
The injectivity of (36) implies that $H^{\bullet}(F_{V})$ vanishes in odd degrees.
Corollary 5.12.
Let $\rho=[r_{1},\ldots,r_{k}]\vdash r$ be a partition of $r$, $V\in M_{\rho}^{\mathrm{ab}}$. Then the weight polynomial (see §4.2) of the fiber $F_{V}$ coincides with its Poincaré polynomial: $E_{t}(F_{V})=P_{t}(F_{V})$.
We are now in a position to prove a crucial recursion determining the Betti numbers of abelian fibers.
@@
Note that $\mathcal{F}_{i}$ is the set of those elements of the fiber $F_{V}$ which have $V_{i}$ as a subbundle. It follows then, that for a subset $J\subset\underline{k}$, the set
@@
Since $\mathcal{F}_{i}$ is a projective bundle (38) with fiber dimension $r_{i}(r-r_{i})(g-1)+r_{i}-1$ over $F_{V/V_{i}}$, its cohomology is pure (cf. Proposition 5.10), and thus we can identify the term in (37) corresponding to $J=\{i\}$ as the weight polynomial of $\mathcal{F}_{i}$.
@@
Now observe that the set $\{\mathcal{F}_{i},\,i=1,\dots,k\}$ is a covering of $F_{V}$, and the corresponding inclusion-exclusion formula results in the identity
@@
This is exactly the equality (37) (cf. Corollary 5.12), and this completes the proof of Theorem 5.13. ∎
@@ 5.4. Intersection of projective towers
Our next project is to study the intersection between two towers, which will be key in §8. Our construction is a generalization of the tower construction in §5.1. To describe this generalization, we will need to introduce a combinatorial device associated with a pair of permutations $\sigma,\tau\in\Sigma_{k}$.
@@
Using this partial order, we construct an ordered decomposition of the set $\underline{k}$
@@
Clearly, this process eventually exhausts the set $\underline{k}$.
@@
    $\mathcal{I}(\sigma,\tau)^{\leq{1}}=\prod_{i\in\Pi_{1}}\mathbb{P}(V_{i|p}).$
@@
    together with the normalized universal bundle (cf. notation on page 3.12 and Lemma 5.2(iii))
@@
Clearly, this happens only if $j<i$. In other words, for any $W\in\mathcal{T}_{\sigma}$ we obtain that $j\prec_{W}i$ implies $\sigma^{-1}(j)<\sigma^{-1}(i)$. Hence, if $W\in\mathcal{T}_{\sigma}\cap\mathcal{T}_{\tau}$, then the partial order $\prec_{W}$ has to be compatible with the partial order
@@
Remark 5.15.
Proposition 5.14 and its proof may be generalized to describe the intersection of any $l$ towers $\bigcap_{\sigma\in I}\mathcal{T}_{\sigma}$ for $I\subseteq\Sigma_{k},\ |I|=l$, by defining the partial order $j\prec i$ as
$j\prec_{I}i\quad\text{if}\quad\sigma^{-1}(j)<\sigma^{-1}(i)\text{ for all }\sigma\in I.$
@@ 5.5. The cohomology ring of the abelian fiber
Our goal is to prove that $\phi$ is an isomorphism. In order to show this, we use the following statement, which we will prove in §7.4.
@@
To prove the statement, we need to show that $\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\subset I$.
Let $d(\rho)=\sum_{1\leq i<j\leq k}r_{i}r_{j}(g-1)$; we start with proving that the surjective morphism of graded rings $\eta$:
@@
induces an isomorphism between the homogeneous degree $2d(\rho)+2r-2k$ elements. It follows from Proposition 5.19 and the recursive formula (37), that the ring $R$ has $k!$ elements of top degree $2d(\rho)+2r-2k$. Since the top-dimensional cohomology of an equidimensional union of projective varieties equals the direct sum of the pieces, we have (see Propositions 5.1 and 5.9(i))
@@
Then, using Lemma 5.17 and the integral formula (34), we obtain that for every $\sigma\in\Sigma_{k}$ the class
@@
under the $\sigma$-th coordinate projection in the direct sum (43). Since the morphism $\theta$ is injective (see the proof of Corollary 5.12), and the $k!$-dimensional space
@@
Next, we repeat the argument from [18, proof of Theorem 7.1]: we note that the ring $R$ can be interpreted as the cohomology ring of the Lawrence toric variety of dimension $d(\rho)+r-k$ (cf. [36, Theorem 6.3] and [18, Theorem 8.3]), and thus the socle of $R$ consists precisely of the elements of top degree $2d(\rho)+2r-2k$. We assume that $\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\not\subset I$ and pick a non-zero element $h\in\cap_{\sigma\in\Sigma_{k}}I_{\sigma}\subset R$ of maximal degree. Our goal is to show that $h$ is zero in $R$, and thus $h\in I$.
@@ 6. The no-support theorem for non-abelian strata
6.1. The cohomological restriction map
Now we turn to the study of non-abelian case. Recall from §3.1 that a complete stratification $\mathcal{M}_{0}(r)=\sqcup_{\overline{\rho}}M_{\overline{\rho}}$ is parametrized by multi-partitions:
@@ 6.1. The cohomological restriction map
A non-abelian stratum $M_{\overline{\rho}}$ carries no support of $R\pi_{*}\mathbb{Q}_{\mathcal{P}_{0}(r)}$, i.e. the corresponding local system $\mathcal{L}_{\overline{\rho}}$ vanishes.
@@
In fact, we will prove a stronger statement. Recall from §3.1 that to the multi-partition $\overline{\rho}$ corresponds a true partition $\rho$ with $k=m_{1}+\dots+m_{s}$ elements. Now let $V\in M_{\rho}^{\mathrm{ab}}\cap\mathcal{C}^{\circ}$; we consider the composition
$H^{\bullet}(F_{\overline{V}})\xrightarrow{j^{*}}H^{\bullet}(\pi^{-1}(\mathcal{C}^{\circ}))\rightarrow H^{\bullet}(F_{V})$
where all maps are pullbacks of the obvious inclusions. Note that the injectivity of the natural restriction map $H^{\bullet}(F_{\overline{V}})\longrightarrow H^{\bullet}(F_{V})$ clearly implies the injectivity of $j^{*}$.
Further, to the multi-partition ${\overline{\rho}}$, we associate the subgroup
$\Sigma[\overline{\rho}]\simeq\Sigma_{m_{1}}\times\dots\times\Sigma_{m_{s}}\subset\Sigma_{k},$
which exchanges the elements of ${\underline{k}}=\{1,2,\dots k\}$. The group $\Sigma[\overline{\rho}]$ acts naturally on the cohomology of the abelian fiber $F_{V}$. The following result is a strengthening of the injectivity of $j^{*}$.
@@
Let $\overline{\rho}$ be a multi-partition and $\rho$ the associated partition, $\overline{V}\in M_{\overline{\rho}}$, and let $V\in M_{\rho}^{\mathrm{ab}}\cap\mathcal{C}^{\circ}$ be as above. The map
@@
The rest of this Section is devoted to the proof of this statement.
6.2. Injectivity and the cohomological restriction map
As in the abelian case, a key simplifying point in our arguments is the purity of the fiber cohomology $H^{\bullet}(\overline{V})$.
Lemma 6.3.
The cohomology of $F_{\overline{V}}$ is pure.
@@ 6.2. Injectivity and the cohomological restriction map
The argument is almost identical to the one given in the abelian case in the proof of Proposition 5.10. While this is not crucial, we note that we can include in the inductive hypothesis that the non-abelian strata are not support strata for the map $\pi$, and thus it is sufficient to study the abelian strata only. The local monodromy is not trivial in this case, and thus instead of part (1) of Proposition 4.6, we use part (2). The map $g_{\rho}$ defined in Remark 3.4 satisfies all conditions on the finite map $g_{\alpha}$ from part (2), and we can again appeal to Proposition 4.6, and Corollary 4.3 to conclude that the cohomology of $F_{\overline{V}}$ is pure. ∎
In the remaining part of this paragraph, we prove that the map (44) is injective.
Proof of injectivity in (44).
To simplify our notation, we will describe in detail the case of $\overline{\rho}=[(s,k)]$, that is, the stratum $M_{\overline{\rho}}$ of bundles $\overline{V}=W^{\oplus k}$, where $W$ is a stable bundle of rank $s$ and degree 0. The general case is obtained by tensoring copies of this basic case (see Remark 6.12).
We thus have $sk=r$, and $\rho=[s,s,\dots,s]$. To avoid double indices, we introduce the notation $lW$ for $W^{\oplus l}$.
Recall (cf. Remark 3.4) that the stratum $M_{\rho}^{\mathrm{ab}}$ may be obtained as a quotient of the space of sequences of stable bundles $\mathcal{M}_{\rho}^{\mathrm{ab}}=\{(V_{1},\dots,V_{k})|\,V_{1}\oplus...\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}\}$ by the symmetric group $\Sigma_{k}$. Then the cohomologies of the abelian fibers $F_{\oplus V_{i}}$ form a trivial bundle over $\mathcal{M}_{\rho}^{\mathrm{ab}}$, to which the action of $\Sigma_{k}$ lifts as the permutation action on the variables $x_{j}$, $j=1,\dots,k$. It follows that the image of the composed map (44) is contained in the invariant part of the cohomology of $H^{\bullet}(F_{V})$; we show now that the map
@@
is injective. Using induction, we assume that the statement holds for $k$ replaced by $k-1$, and our goal is to prove it for $k$.
@@
where the map $f^{\prime}$ is a smooth fibration. We introduce the notation
@@
for the composed map $\widehat{\pi}=\pi\circ h^{\prime}$. In fact, we will be mainly interested in the subset of $\hat{\mathcal{E}}$ lying over the closure of $M_{\rho}^{\mathrm{ab}}\subset\mathcal{M}_{0}(ks)$.
It follows from the proof of Proposition 5.1 that for an abelian $V=V_{1}\oplus\dots\oplus V_{k}\in M_{\rho}^{\mathrm{ab}}$ we have
@@
The purity of $H^{\bullet}(F_{V})$ (see Proposition 4.6) and that of the cohomology of the sets $\mathcal{F}_{j}$ (see (38)), which form a covering of $F_{V}$, imply that we have a $\Sigma_{k}$-invariant embedding $H^{\bullet}(F_{V})\hookrightarrow\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$, and this provides us with an embedding
(48) ${\bar{Q}}:H^{\bullet}(F_{V})^{\Sigma_{k}}\hookrightarrow\left[\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})\right]^{\Sigma_{k}}.$
@@
By Lemma 6.3, $H^{\bullet}(F_{kW})$ has pure Hodge structure, and so does $H^{\bullet}(\mathcal{E})$, since $\mathcal{E}$ is a bundle of projective spaces over $F_{(k-1)W}$. Now the injectivity follows from Proposition 4.1, and the inductive hypothesis. ∎
Finally, we note that the cohomology $H^{\bullet}(\widehat{\pi}^{-1}(V))$ of the fibers (47) over $M_{\rho}^{\mathrm{ab}}$ may be obtained as a descent to $M_{\rho}^{\mathrm{ab}}$ of the trivial bundle $\oplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$ on $\mathcal{M}_{\rho}^{\mathrm{ab}}$ endowed with a $\Sigma_{k}$-action. This results in a natural map
@@
Using descent with respect to the subgroup $\Sigma_{k-1}\subset\Sigma_{k}$, we can conclude that the projection of the map $H^{\bullet}(\mathcal{E})\longrightarrow\bigoplus_{j=1}^{k}H^{\bullet}(\mathcal{F}_{j})$ to $H^{\bullet}(\mathcal{F}_{1})$ has its image in $H^{\bullet}(\mathcal{F}_{1})^{\Sigma_{k-1}}$, and we recall that, by our inductive hypothesis,
@@
Clearly, Lemma 6.5 and the commutative diagram (50) imply that the map (44) is injective. ∎
This already proves Theorem 6.1. For completeness, however, we will show below that this map is an isomorphism.